Two identical objects of mass m are a distance R apart and feel a gravitational force of attraction of F.

If the distance increases to 5R, what will be the force between the two objects?

Select one:

a. F/5

b. F/25

c. F

d. 25F

e. 5F
Answer #1

here,

when Two identical objects of mass m are a distance R apart

the gravitational force , F = G * m^2 /R^2

when the distance is 5R

the new gravitational force , F' = G * m^2 /(5R)^2

F' = 1/25 *(G * m^2 /R^2)

F' = F/25

the new gravitational force is b) (F/25)